yeetantor:

A dance school has 63 ballet dancers. If 42% of the total number of dancers in the school are ballet dancers, what is the total number of dancers in the school? sorry need help again

yeetantor:

hi again

jimthompson5910:

How far did you get with this one?

yeetantor:

not really far at all im kinda stupid when it comes to math sooooooooo

jimthompson5910:

so what I like to do when solving problems is to jump immediately to the question at hand "what is the total number of dancers in the school?" We don't know this value, so lets call it x x = number of dancers at the school

yeetantor:

ok

jimthompson5910:

we are told that 42% of all the dancers are ballet dancers, and there are 63 ballet dancers this means 42% of x = 63 Agreed?

yeetantor:

agreed

jimthompson5910:

change the 42% into 0.42 this is because 42% = 42/100 = 0.42 Also change the "of" into a multiplication symbol We go from 42% of x = 63 to 0.42*x = 63

jimthompson5910:

how can we solve for x in 0.42*x = 63 ?

yeetantor:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 how can we solve for x in 0.42*x = 63? uhhhhhhhhh..............

yeetantor:

idk

jimthompson5910:

we need to undo that multiplication somehow what is the opposite of multiplication?

yeetantor:

um division

jimthompson5910:

yes, so we'll divide both sides by 0.42

jimthompson5910:

On the left side, when you divide by 0.42, the "0.42" terms go away On the right side, what is 63/0.42 equal to?

yeetantor:

umm its equal to i don't know anything about mathhhhhhhhh

jimthompson5910:

use a calculator

yeetantor:

ok hold on what am using a calcultor for

jimthompson5910:

you'll be dividing 63 over 0.42

yeetantor:

ok so i divide 63 to 0.42

jimthompson5910:

yes

yeetantor:

ok hold on

yeetantor:

150

jimthompson5910:

yes

jimthompson5910:

so thats the answer

yeetantor:

ok thanks

jimthompson5910:

sure thing